# Discovery of energy landscapes towards optimized quantum transport: Environmental effects and long-range tunneling

### v1

@@ II.4 Optimization
For OGA [96], in step $k$, the site energies are updated according to
$\vec{\varepsilon}_{k+1}=\vec{\varepsilon}_{k}+2h\nabla_{\varepsilon}\eta\left(\vec{\varepsilon}_{k}\right)-h\nabla_{\varepsilon}\eta\left(\vec{\varepsilon}_{k-1}\right),$ (10)
and in AdaMax [97, 98], the site energies are updated according to
@@
Eq. (12) provides a stopping condition for the gradient ascent algorithm: exit the loop if $\left|\nabla_{\varepsilon}\eta\left(\vec{\varepsilon}_{k}\right)\right|<\epsilon$, where $\epsilon$ is some small tolerance. Python libraries JAX and Optax [97] are used to calculate the flux derivatives and perform the OGA and AdaMax optimization.
@@
While for chains with few sites it is possible to exhaustively optimize over the entire hypergrid, this becomes computationally impractical, and generally unnecessary, for larger systems. In such cases, random initializations followed by gradient ascent optimization provide a sufficient strategy for exploring the landscape of possible energy configurations.
In Appendix A we compare the optimization results for a three-level system using different optimization algorithms. Ultimately, OGA and AdaMax were selected because they performed the best on our problems, which typically included a “ridge” of close-to-optimal solutions.
II.5 Parameters and transport mechanisms
@@
Depending on the range of parameters, in the coherent case one can observe sequential tunneling, deep tunneling, or ballistic motion. Furthermore, nonzero long-range terms, such as $J_{2}$, lead to interference effects with, e.g., short-range tunneling terms, $J_{1}$. Such interference effects can be constructive or destructive (see Section III). When environmental effects are included, they can enhance the flux by destroying the Anderson localization effect through level broadening, or suppress ballistic transport. Strong environmental effects in the form of local dephasing are known to freeze dynamics, interpreted as the quantum Zeno effect [47, 50, 54]. Furthermore, the finite-temperature dissipation model (OQS II) can enact diffusion-like transport [99].
@@
In this study, we select parameters for the chains such that competing transport effects contribute. That is, we set $J_{max}$, $\gamma_{l}$, the dephasing rate constant $\Gamma$ and the temperature $T$ to have similar values, thus achieving transport where they all compete and contribute.
@@
Figure 2: Coherent Model: population flux in the absence of environmental effects. Left, (a)-(c): 3-site model with nearest-neighbor tunneling only, $J_{1}=0.2$ and $J_{2}=0$. (a) Flux map as a function of $\varepsilon_{2}$ and $\varepsilon_{3}$, fixing $\varepsilon_{1}=0$. (b) Level diagram of the optimal configuration, $\varepsilon_{2}\in[-3.0,3.0]\cdot 10^{-4}$ (these values yield the same flux) and $\varepsilon_{3}=1.6\cdot 10^{-11}$, with the optimal flux $\eta=0.032$. (c) Steady-state population under the optimal profile, $\rho_{11}^{SS}=\rho_{22}^{SS}=0.34$, $\rho_{33}^{SS}=0.32$. Right, (d)-(e): 3-site model with beyond-nearest-neighbor tunneling, $J_{1}=0.2$ and $J_{2}=0.1$. (d) Population flux map. (e) Level diagram of the optimal configuration, $\varepsilon_{2}=-0.200$ and $\varepsilon_{3}=-0.050$ leading to $\eta=0.033$. (f) Steady-state population under the optimal profile, $\rho_{11}^{SS}=0.37$, $\rho_{22}^{SS}=0.30$, $\rho_{33}^{SS}=0.33$ Other parameters are $\Gamma=0$, leakage rate constant $\gamma_{l}=0.1$. In panels (a) and (d), results of the OGA algorithm are marked by a black x.
Figure 3: OQS Model I: population flux in the presence of environmental effects with $\Gamma=0.1$. Left, (a)-(c): model with nearest-neighbor tunneling, $J_{1}=0.2$ and $J_{2}=0$. (a) Population flux map. (b) Level diagram of the optimal configuration, $\varepsilon_{2}=1.3\cdot 10^{-5}$, $\varepsilon_{3}=5.8\cdot 10^{-6}$, resulting in $\eta=0.029$. (c) Steady-state population under the optimal profile, $\rho_{11}^{SS}=0.37$, $\rho_{22}^{SS}=0.34$, $\rho_{33}^{SS}=0.29$. Right, (d)-(e): 3-site model next-nearest-neighbor tunneling, $J_{1}=0.2$ and $J_{2}=0.1$. (d) Population flux map. (e) Level diagram of the optimal configuration, $\varepsilon_{2}=-0.292$, $\varepsilon_{3}=-0.017$, leading to $\eta=0.030$. (f) Steady-state population under the optimal profile, $\rho_{11}^{SS}=0.36$, $\rho_{22}^{SS}=0.34$, $\rho_{33}^{SS}=0.30$. Other parameters are $\gamma_{l}=0.1$. In both (a) and (d), optimization results of the OGA algorithm are marked by a black x.
@@ III Optimal Energy Landscapes in three-site systems
We begin our analysis by testing the optimization procedure on a minimal model, a three-site system. In this setup, the population leaves site 3 at a fixed rate constant $\gamma_{l}$ and is injected back to site 1 at the same rate constant, achieving a steady state. Without loss of generality, we set the first site energy, $\varepsilon_{1}$, to zero. We also fix the tunneling elements. The parameters varied in our model are the site energies $\varepsilon_{2}$ and $\varepsilon_{3}$; by tuning these, we seek to optimize the population flux defined in Eq. (9). Since only two parameters are varied in the optimization process, we can visualize the flux landscape as a contour map. We also easily compare the results of the optimization algorithm with the true global and local maxima obtained via extensive brute-force simulations.
@@ III.1 Coherent model
We begin with the coherent case, obtained by setting $\Gamma_{n}=0$ in Eq. (5). Fig. 2 shows simulation results for the three-level system under two scenarios: a model with nearest-neighbor interactions only (left), and a model with an additional tunneling element between the first and third levels (right). In both scenarios, we present a complete flux map as a function of the two energy parameters, $\varepsilon_{2}$ and $\varepsilon_{3}$ while keeping $\varepsilon_{1}=0$. The point marked with an “x” indicates the value obtained via the OGA optimization protocol; it clearly successfully identified the optimal energy landscape for the system.
From Fig. 2, we draw the following observations: (i) When only nearest-neighbor coupling is allowed, i.e. $J_{1}\neq 0$ and $J_{2}=0$ (left panels), the optimal energy landscape is nearly flat with $\varepsilon_{2,3}\approx\varepsilon_{1}$; see Fig. 2(a)-(b). Notably, a broad range of values for the intermediate level $\varepsilon_{2}$ supports close to optimal transfer fluxes. In contrast, when the next-nearest-neighbor coupling is introduced, i.e. $J_{1}\neq 0$, $J_{2}\neq 0$, a V-shaped energy profile leads to enhanced transport and is optimal; see Fig. 2(d)-(e). In this structure, the middle site (site 2) is essentially moved “out of the way”, since the transfer can be facilitated by a direct tunneling transition from site 1 to site 3. Another notable observation is that the optimal configuration shows populations close to $1/3$. We recall that optimal flux corresponds to high population at the exit site (3), which in this system corresponds to a value approaching 1/3.
@@
This simple result captures rich behavior: assuming that $\gamma_{l}>J_{1}$, that is, that the leakage rate is not the slower, rate-determining step, we can distinguish between the deep tunneling limit, $\varepsilon_{2}\gg J_{1}$, and the ballistic limit, $\varepsilon_{2}\ll J_{1}$,
$\eta(J_{2}=0,\Gamma=0)=\begin{cases}\frac{4J_{1}^{4}}{\gamma_{l}\varepsilon_{2}^{2}}&{\rm for\,\,}\varepsilon_{2}\gg J_{1}\\ \frac{2J_{1}^{2}}{\gamma_{l}}&{\rm for\,\,}\varepsilon_{2}\ll J_{1}.\end{cases}$ (14)
In the opposite limit of a small leakage rate, $\gamma_{l}\ll\varepsilon_{2}$ and $\gamma_{l}\ll J_{1}$, we find $\eta=\gamma_{l}/3$, reflecting that the system reaches an infinite-temperature equal-population state, and that the flux is dictated by the leakage rate, which is the slowest, thus rate-determining, step. However, in our simulations, the parameters $\gamma_{l}$ and $J_{1}$ are of comparable magnitude. As such, neither of the limiting cases discussed above fully captures the observed behavior.
We now examine the role of next-nearest-neighbor tunneling, $J_{2}\neq 0$. The introduction of this tunneling element opens the door to interference effects. As before, we simplify the analysis by setting $\varepsilon_{1}=\varepsilon_{3}=0$, and neglecting environmental effects. In this case, the population flux is
@@
The analysis of the coherent case in Fig. 2 provides three key guidelines for optimizing transport: (i) In the absence of environmental effects there are many solutions (energy profiles) that are near optimal. These solutions cover a wide range of values for the intermediate level, $\varepsilon_{2}$. (ii) When multiple pathways are allowed due to next-nearest-neighbor tunneling, destructive interference effects appear when all energies are near resonance. To suppress these interference effects and achieve high flux, the intermediate level needs to be detuned from the entry and exit sites. That is, $|\varepsilon_{2}|\gg|\varepsilon_{1}|$, $|\varepsilon_{3}|$ promotes transport. (iii) The existence of multiple pathways, together with the openness of the system due to leakage, $\gamma_{l}\neq 0$, breaks the symmetry of the energy profile with respect to $\varepsilon_{2}$, see Fig. 2(d).
@@ III.2 OQS Model I
We now incorporate environmental effects into the chain, in the form of local dephasing as described in Sec. II.1. Figure 3 presents results for the three-level system, comparing the nearest-neighbor tunneling case [panels (a)–(c)] with the scenario that includes next-nearest-neighbor tunneling [panels (d)–(f)]. When only nearest-neighbor tunnelings are allowed (left), the flux depends on the interplay of coherent and incoherent effects. With the present parameters, this still favors a nearly flat energy profile, similar to the $\Gamma=0$ case of Fig. 2(a)-(b). When next-nearest-neighbor tunneling is included (right), interference effects take place, again similarly to the $\Gamma=0$ case, with the flux optimized near $\varepsilon_{2}\approx-0.3$. Comparing Fig. 2 ($\Gamma=0$) with Fig. 3 ($\Gamma=0.1$), we observe that transfer trends are similar, although in the latter case an asymmetry with respect to the energy $\varepsilon_{3}$ develops.
@@
$+$ $\Gamma\frac{2\gamma_{l}^{2}(\gamma_{l}^{2}\varepsilon_{2}^{2}J_{1}^{2}-8\varepsilon_{2}^{2}J_{1}^{4}-12J_{1}^{6})}{(\gamma_{l}^{2}\varepsilon_{2}^{2}+2\gamma_{l}^{2}J_{1}^{2}+12J_{1}^{4})^{2}}+O(\Gamma^{2})$
Once again, we distinguish between the deep tunneling limit ($\varepsilon_{2}\gg J_{1}$) and the ballistic regime ($\varepsilon_{2}\ll J_{1}$), resulting in
@@
As expected, environmental effects assist transport in the deep tunneling regime, but reduce the flux when transfer occurs ballistically.
@@
From the second term, we note that the effect of dephasing is to suppress the short-range ($J_{1}$) ballistic transport. The last term is proportional to $J_{2}^{2}\Gamma$. The sign of this last term depends on the relative magnitude of $\gamma_{l}$ and $J_{1}$. As such, environmental effects can either enhance or suppress the long-range $J_{2}$ transport channel. Specifically, when $\gamma_{l}\ll J_{1}$ this last term becomes negative, reflecting destructive interference that reduces the flux. However, it is important to remember that Eq. (19) assumes a flat energy profile, and as such does not capture the full complex picture that we present in Fig. 3 when $J_{2}\neq 0$.
This discussion is presented here to provide an appreciation of the range of transport regimes that appear in a system as short as three sites.

### v2

@@ II.4 Optimization
At the heart of this work lies a fundamental question: which classes of energy profiles, $\vec{\varepsilon}=(\varepsilon_{1},\varepsilon_{2},...,\varepsilon_{N})$, maximize transport flux? We focus in particular on how long-range versus short-range tunneling shapes this optimization, and how environmental coupling and finite-temperature effects further influence the optimal energy landscape. Because the dynamics entwine both coherent and incoherent processes, the resulting optimal profiles are expected to emerge from the interplay between these competing mechanisms.
@@
For OGA [76], in step $k$, site energies are updated according to
$\vec{\varepsilon}_{k+1}=\vec{\varepsilon}_{k}+2h\nabla_{\varepsilon}\eta\left(\vec{\varepsilon}_{k}\right)-h\nabla_{\varepsilon}\eta\left(\vec{\varepsilon}_{k-1}\right).$ (10)
In AdaMax [27, 57], site energies are updated according to
@@
Eq. (12) provides a stopping condition for the gradient ascent algorithm: exit the loop if $\left|\nabla_{\varepsilon}\eta\left(\vec{\varepsilon}_{k}\right)\right|<\epsilon$, where $\epsilon$ is some small tolerance. No constraints were placed on the energies – instead, a maximum number of steps was defined to stop diverging optimizations from iterating indefinitely. Python libraries JAX and Optax [27] are used to calculate the flux derivatives and perform the OGA and AdaMax optimization.
@@
For chains with only a few sites, an exhaustive search over the full hypergrid is feasible. However, as system size grows, this approach quickly becomes computationally prohibitive, and, in practice, unnecessary. In larger systems, efficient exploration of the energy landscape can be achieved through random initializations followed by gradient ascent optimization.
In Appendix A, we benchmark several optimization algorithms on a three-level system. Among them, OGA and AdaMax consistently performed best for our class of problems, which often feature a characteristic “ridge” of near-optimal solutions.
II.5 Discussion over transport mechanisms and choice of parameters
@@
Depending on the range of parameters, in the coherent case one can observe sequential tunneling, deep tunneling, or ballistic motion. Furthermore, long-range tunneling coupling terms, such as $J_{2}$, lead to interference effects with, e.g., short-range processes emerging from tunneling terms $J_{1}$. Such interference effects can be constructive or destructive (see Section III). When environmental effects are included, they can enhance the flux by destroying the Anderson localization effect through level broadening, or suppress ballistic transport. Strong environmental effects in the form of local dephasing are known to freeze dynamics, interpreted as the quantum Zeno effect [85, 115, 70]. Furthermore, the finite-temperature dissipation model (OQS II) can enact diffusion-like transport [65].
We now provide a more organized discussion of transport regimes in quasi one-dimensional systems. For a detailed discussion, we refer readers to reviews such as [30, 62]:
(i) Quantum coherent transport: when the dephasing rate is zero or much smaller than the tunneling amplitudes ($\Gamma_{n}\ll J_{|n-m|}$), the dynamics and steady state are governed by the unitary (coherent) part of the evolution. We therefore refer to this regime as coherent transport. Within this regime, several distinct scenarios can be identified depending on the system’s energy landscape. If the terminal sites lie significantly lower or higher in energy than the intermediate sites, deep tunneling processes occur, characterized by an exponentially decaying flux with increasing system length. In contrast, for a flat energy profile, ballistic transport arises, where transport measures such as flux, current, or efficiency become independent of system size. Disordered systems, on the other hand, are expected to exhibit localization in the thermodynamic limit. Since we do not analyze the scaling of transport properties with system length in this work, we infer possible underlying transport mechanisms directly from steady-state features, the coherences and population patterns. A uniform (flat) population profile, for example, is indicative of ballistic motion.
(ii) Incoherent regime: when the dephasing or dissipation rate dominates all other relevant energy scales, i.e., the tunneling elements, site energies, and leakage rates ($\Gamma\gg J_{|n-m|},\varepsilon_{n},\gamma_{l}$), transport becomes diffusive with system length. In this regime, transport measures (e.g., flux or current) are expected to scale as $1/N$. Since we do not explicitly analyze the length dependence here, we instead use Fick’s law, manifested as a constant population gradient, as an indicator of diffusive motion. Furthermore, the dependence of the flux on $\Gamma$ manifests a Kramer’s turning behavior from $\eta\propto\Gamma$ in the weak dissipative limit to $\eta\propto 1/\Gamma$ under strong dissipation [111].
(iii) Intermediate regime: when the dephasing rate $\Gamma$ and the dissipation rate are of a comparable magnitude to the tunneling amplitudes, $J_{|n-m|}$, the dynamics and resulting steady state reflect a competition between coherent and incoherent processes. In this regime, neither $\Gamma$ nor the tunneling terms can be neglected (see, for example, Eq. (16)), and the resulting flux arises from the influence of both effects. This interplay can, in some cases, enhance transport efficiency, particularly in energetically disordered systems, a phenomenon known as environmental noise-assisted quantum transport.
@@
In this study, we select parameters for the chains such that competing transport effects contribute to the population flux. That is, we set $J_{max}$, $\gamma_{l}$, the dephasing rate constant $\Gamma$ and the temperature $T$ to have similar values, thus achieving transport where they all participate. The purpose of the study was not to identify dominant transport mechanisms; this discussion is meant to highlight different possible underlying transport mechanisms that dominate in different regimes.
@@
As for the magnitude of the population flux, given the definition, Eq. (9), for an $N$-site chain we anticipate it to be about $\eta\approx\gamma_{l}/N$ in the steady state limit (assuming for simplicity equal population). As such, for our choice of parameters and for $N=10$ sites, we get $\eta=0.01$, which translates to $\eta\approx$ 16 ps-1 if we consider parameters in units of eV.
Figure 2: Coherent Model: population flux in the absence of environmental effects. Left, (a)-(c): 3-site model with nearest-neighbor tunneling, $J_{1}=0.2$ and $J_{2}=0$. (a) Flux map as a function of $\varepsilon_{2}$ and $\varepsilon_{3}$, fixing $\varepsilon_{1}=0$. (b) Level diagram of the optimal configuration, $\varepsilon_{2}\in[-3.0,3.0]\cdot 10^{-4}$ (these values yield the same flux) and $\varepsilon_{3}=1.6\cdot 10^{-11}$, with the optimal flux $\eta=0.032$. (c) Steady-state population under the optimal profile, $\rho_{11}^{SS}=\rho_{22}^{SS}=0.34$, $\rho_{33}^{SS}=0.32$. Right, (d)-(e): 3-site model with beyond-nearest-neighbor tunneling, $J_{1}=0.2$ and $J_{2}=0.1$. (d) Population flux map. (e) Level diagram of the optimal configuration, $\varepsilon_{2}=-0.200$ and $\varepsilon_{3}=-0.050$ leading to $\eta=0.033$. (f) Steady-state population under the optimal profile, $\rho_{11}^{SS}=0.37$, $\rho_{22}^{SS}=0.30$, $\rho_{33}^{SS}=0.33$. Other parameters are $\Gamma=0$, leakage rate constant $\gamma_{l}=0.1$. In panels (a) and (d), results of the OGA algorithm are marked by a black x.
Figure 3: OQS Model I: population flux in the presence of environmental effects with $\Gamma=0.1$. Left, (a)-(c): model with nearest-neighbor tunneling, $J_{1}=0.2$ and $J_{2}=0$. (a) Population flux map. (b) Level diagram of the optimal configuration, $\varepsilon_{2}=1.3\cdot 10^{-5}$, $\varepsilon_{3}=5.8\cdot 10^{-6}$, resulting in $\eta=0.029$. (c) Steady-state population under the optimal profile, $\rho_{11}^{SS}=0.37$, $\rho_{22}^{SS}=0.34$, $\rho_{33}^{SS}=0.29$. Right, (d)-(e): 3-site model with next-nearest-neighbor tunneling, $J_{1}=0.2$ and $J_{2}=0.1$. (d) Population flux map. (e) Level diagram of the optimal configuration, $\varepsilon_{2}=-0.292$, $\varepsilon_{3}=-0.017$, leading to $\eta=0.030$. (f) Steady-state population under the optimal profile, $\rho_{11}^{SS}=0.36$, $\rho_{22}^{SS}=0.34$, $\rho_{33}^{SS}=0.30$. We set $\gamma_{l}=0.1$. In both (a) and (d), optimization results of the OGA algorithm are marked by a black x.
@@ III Optimal Energy Landscapes in three-site systems
We begin by testing the optimization procedure on a minimal three-site model. Detailed analysis of such models were carried out in, e.g., Refs. [11, 67], with a particular focus on multi-path destructive interference processes. In this setup, population exits from site 3 at a fixed rate constant $\gamma_{l}$ and is injected back to site 1 at the same rate constant, establishing a steady state. Without loss of generality, we set the first site energy, $\varepsilon_{1}$, to zero. We also fix the tunneling elements. The remaining site energies, $\varepsilon_{2}$ and $\varepsilon_{3}$, serve as the optimization variables, with the objective of maximizing the population flux defined in Eq. (9). Since only two parameters are varied in the optimization process, we can visualize the flux landscape as a contour map. We can also easily compare the results of the optimization algorithm with the true global and local maxima obtained via extensive brute-force simulations.
@@ III.1 Coherent model
We begin with the coherent case, obtained by setting $\Gamma_{n}=0$ in Eq. (5). Fig. 2 shows simulation results for the three-level system under two scenarios: a model with nearest-neighbor interactions only (left), and a model with an additional tunneling element between the first and third levels (right). In both scenarios, we present a complete flux map as a function of the two energy parameters, $\varepsilon_{2}$ and $\varepsilon_{3}$ while keeping $\varepsilon_{1}=0$. The point marked with an “x” indicates the scenario obtained via the OGA optimization protocol; it clearly successfully identified the optimal energy landscape for the system.
From Fig. 2, we draw the following conclusions: (i) When only nearest-neighbor coupling is allowed, i.e. $J_{1}\neq 0$ and $J_{2}=0$ (left panels), the optimal energy landscape is nearly flat with $\varepsilon_{2,3}\approx\varepsilon_{1}$; see Fig. 2(a)-(b). Notably, a broad range of values for the intermediate level $\varepsilon_{2}$ supports close to optimal transfer fluxes. In contrast, when the next-nearest-neighbor coupling is introduced, i.e. $J_{1}\neq 0$, $J_{2}\neq 0$, a V-shaped energy profile leads to enhanced transport and is optimal; see Fig. 2(d)-(e). In this structure, the middle site (site 2) is essentially moved “out of the way”, since the transfer can be facilitated by a direct tunneling transition from site 1 to site 3. Another notable observation is that optimal configurations show populations close to $1/3$. We recall that optimal flux corresponds to high population at the exit site (3), which in this system corresponds to a value approaching 1/3.
@@
This simple result captures rich behavior: assuming that $\gamma_{l}\gg J_{1}$, that is, the leakage rate is not the slower, rate-determining step, we can distinguish between the deep tunneling limit, $\varepsilon_{2}\gg J_{1}$, and the case $\varepsilon_{2}\ll J_{1}$, which would evolve to band motion – ballistic dynamics in long chains,
$\eta(J_{2}=0,\Gamma=0;\gamma_{l}\gg J_{1})=\begin{cases}\frac{4J_{1}^{4}}{\gamma_{l}\varepsilon_{2}^{2}}&{\rm for\,\,}\varepsilon_{2}\gg J_{1}\\ \frac{2J_{1}^{2}}{\gamma_{l}}&{\rm for\,\,}\varepsilon_{2}\ll J_{1}.\end{cases}$ (14)
In the opposite limit of a small leakage rate, $\gamma_{l}\ll\varepsilon_{2}$ and $\gamma_{l}\ll J_{1}$, we find $\eta=\gamma_{l}/3$, reflecting that the system reaches an equal-population state, and that the flux is dictated by the leakage rate, which is the slowest, thus rate-determining, step. However, in our simulations, the parameters $\gamma_{l}$ and $J_{1}$ are of comparable magnitude. As such, neither of the limiting cases discussed above fully captures the observed behavior.
We now analyze analytically the role of next-nearest-neighbor tunneling, $J_{2}\neq 0$. The introduction of this tunneling element opens the door to interference effects. As before, we simplify the analysis by setting $\varepsilon_{1}=\varepsilon_{3}=0$, and neglecting environmental effects. In this case, the population flux is
@@
The analysis of the coherent case in Fig. 2 provides three key guidelines for optimizing transport: (i) In the absence of environmental effects there are many solutions (energy profiles) that are near optimal. These solutions cover a wide range of values for the intermediate level, $\varepsilon_{2}$. (ii) When multiple pathways are allowed due to next-nearest-neighbor tunneling, destructive interference effects appear when non-neighboring sites’ energies are near resonance. To suppress these interference effects and achieve high flux, the intermediate level needs to be detuned from the entry and exit sites. That is, from Eq. (15) we see that, under the assumption of $J_{2}\ll J_{1}$, either setting $\varepsilon_{2}<0$ or $\varepsilon_{2}>J_{1}^{2}/J_{2}$ promotes transport. (iii) The existence of multiple pathways, together with the openness of the system due to leakage, $\gamma_{l}\neq 0$, breaks the symmetry of the energy profile with respect to $\varepsilon_{2}$, see Fig. 2(d).
@@ III.2 OQS Model I
We now incorporate environmental effects into the chain, in the form of local dephasing as described in Sec. II.1. Figure 3 presents results for the three-level system, comparing the nearest-neighbor tunneling case [panels (a)–(c)] with the scenario that includes next-nearest-neighbor tunneling [panels (d)–(f)]. When only nearest-neighbor tunnelings are allowed (left), the flux depends on the interplay of coherent and incoherent effects. With the present parameters, this still favors a nearly flat energy profile, similar to the $\Gamma=0$ case of Fig. 2(a)-(b). When next-nearest-neighbor tunneling is included (right), interference effects take place, again similarly to the $\Gamma=0$ case, with the flux optimized near $\varepsilon_{2}\approx-0.3$. Comparing Fig. 2 ($\Gamma=0$) with Fig. 3 ($\Gamma=0.1$), we observe that in such short systems, transfer trends are similar, although in the latter case an asymmetry with respect to the energy $\varepsilon_{3}$ develops.
@@
$+$ $\frac{2\gamma_{l}^{2}(\gamma_{l}^{2}\varepsilon_{2}^{2}J_{1}^{2}-8\varepsilon_{2}^{2}J_{1}^{4}-12J_{1}^{6})}{(\gamma_{l}^{2}\varepsilon_{2}^{2}+2\gamma_{l}^{2}J_{1}^{2}+12J_{1}^{4})^{2}}\Gamma+O(\Gamma^{2})$
Once again, we distinguish between the deep tunneling limit ($\varepsilon_{2}\gg J_{1}$) and the near-resonant regime ($\varepsilon_{2}\ll J_{1}$), resulting in
@@
As expected, environmental effects assist transport in the deep tunneling regime, but reduce the flux when transfer occurs near resonance.
In the opposite limit of large dephasing compared to the coherent energy parameters, $\Gamma\gg\varepsilon_{2},J_{1},\gamma_{l}$ we get from Eq. (16)
$\eta(J_{2}=0)\approx\frac{J_{1}^{2}}{6\Gamma}.$ (19)
The flux is inversely proportional to $\Gamma$, typical to the quantum Zeno effect or the (classical) large friction limit [45]
@@
From the second term, we note that the interplay of dephasing and short-range ($J_{1}$) coherent tunneling is to suppress transport. The last term is proportional to $J_{2}^{2}\Gamma$. The sign of this term depends on the relative magnitude of $\gamma_{l}$ and $J_{1}$. As such, the interaction of environmental effects with the long-range $J_{2}$ coupling is either to enhance or suppress transport. Specifically, when $\gamma_{l}\ll J_{1}$ this last term becomes negative, reflecting destructive interference that reduces the flux. However, it is important to remember that Eq. (20) assumes a flat energy profile, and as such does not capture the full complex picture that we present in Fig. 3 when $J_{2}\neq 0$. This discussion is presented here to provide an appreciation of the range of transport regimes that appear in a system as short as three sites.
Finally, we point out that, as expected, Eqs. (13)-(20) reveal that the flux depends on the leak rate, $\gamma_{l}$. This dependence is physical: the leak process represents an experimental extraction parameter that controls the rate at which population is collected from the exit site.